Convert the following Hexadecimal values to their Decimal equivalents:
33

9C

FD

12

40
Answer #1

1. 51

  • 33 = (3*161) + (3*160)

2. 156

  • 9C = (9*161) + (12*160)

3. 253

  • FD = (15*161) + (13*160)

4. 18

  • 12 = (1*160) + (2*160)

5. 64

  • 40 = (4*160) + (0*161)

Hence, these are the decimal equivalents.